Use the contingency table to complete A) through D)

EVENT A EVENT B
EVENT C 29 29
EVENT D 2 15
EVENT E 24 1

DETERMINE THE probability

A) P(A/C)=

B) P(C/A)=

C) P(B/E)=

D) P(E/B)=
Answer #1

a)

P(A | C) = P( A and C) / P(C)

= 29 / (29 + 29)

= 0.5

b)

P( C | A) = P( C and A) / P(A)

= 29 / ( 29 + 2 + 24)

= 0.5273

c)

P(B | E) = P( B and E) / P(E)

= 1 / ( 24 + 1)

= 0.04

d)

P(E | B) = P( E and B) / P(B)

= 1 / ( 29 + 15 + 1)

= 0.0222